Answer ALL problems using the bubble sheet given(w
Answer #1

Charge of 1 electron = 1.6*10^-19 C
Net Charge Q, = 1 C

No of Electrons, =  1/(1.6*10^-19)
N = 6.3 * 10^18

2)
Electrical Field, E = k*q/r^2
E = 8.9*10^9 * q^2)/
Force = m*a
Electrical Force = k*q^2/r^2
k*q^2/r^2 = m*a
(8.9*10^9*q^2)/(6.4/100)^2 = 3.2*10^-3 * 538
q = 8.9*10^-7
q = 890 nC

Please post seperate questions in seperate post only.